A ball is thrown vertically upward with a speed of 15 m/s.


A. what is its maximum height

B. when will it return

C. where will it be after 1 s ?


1
Expert's answer
2022-02-10T13:37:59-0500

Given:

"v_0=\rm15\: m/s"

"g=9.8\:\rm m/s^2"


A. the maximum height

"h_{\max}=\frac{v_0^2}{2g}=\frac{15^2}{2*9.8}=11.5\:\rm m"

B. total time of motion

"t=2*t_{up}=2*\frac{v_0}{g}=2*\frac{15}{9.8}=3.1\:\rm s"

C. at the instant "t=1\:\rm s"

"h=v_0t-gt^2/2\\
=15*1-9.8*1^2/2=10.1\:\rm m"